THE B ^{2}Σ^{−} − X^{2}Π_{3/2} TRANSITION OF AuO

\underline{LEAH C. O’BRIEN}, BRADLEY A. BORCHERT, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN, SADASIVAN SHAJI, ADAM FAR-QUHAR, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499.

The visible spectrum of AuO has been recorded in the 16700 - 18700 cm^{−1}region using Intracavity Laser Spectroscopy (ILS). Numerous bands were observed in the spectrum. We have identified 4 new bands with a similar branch structure and regular spacing as being part of a vibrational progression of a new electronic transition which we have tentatively identified as the B ^{2}Σ^{−} − X^{2}Π_{3/2} transition. Results of the analysis will be presented.